OpenStudy (anonymous):

HELP, I will medal & fan ! A pair of linear equations is shown below: y = -x + 1 y = 2x + 4 Which of the following statements best explains the steps to solve the pair of equations graphically? (4 points)

OpenStudy (anonymous):

On a graph, plot the line y = -x + 1, which has y-intercept = -1 and slope = 1, and y = 2x + 4 which has y-intercept = 2 and slope = 4, and write the coordinates of the point of intersection of the two lines as the solution. On a graph, plot the line y = -x + 1, which has y-intercept = 1 and slope = 1, and y = 2x + 4 which has y-intercept = 1 and slope = 4, and write the coordinates of the point of intersection of the two lines as the solution. On a graph, plot the line y = -x + 1, which has y -intercept = 1 and slope = -1, and y = 2x + 4 which has y-intercept = -2 and slope = 2, and write the coordinates of the point of intersection of the two lines as the solution. On a graph, plot the line y = -x + 1 which, has y-intercept = 1 and slope = -1, and y = 2x + 4 which has y-intercept = 4 and slope = 2, and write the coordinates of the point of intersection of the two lines as the solution.

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

the slope of the first eq is -1

OpenStudy (anonymous):

y=-x+1 the form is y=mx+b m is the slope so the slope is -1 y intercept, when you= 0 y=-x+1 0=-x+1 solve for x x=1

OpenStudy (anonymous):

Thank-you for helping me can you help me with more ?

OpenStudy (anonymous):

@AidenS